OpenStudy (anonymous):

what is the value of x in the following equation? 3(x + 2) = 6x + 9

OpenStudy (anonymous):

3x +6 =6x +9 3x -6x = -6+9 -3x = 3 x=-1

OpenStudy (anonymous):

\[3x + 6=6x +9 => 3x=6-9 => x=-1\]